NAOJ GW Elog Logbook 3.2

[Aritomi, Yuhang]
According to Henning result, the shot noise should be -132dBV for 2mW of incident light. In our case, the machine noise is -100dBV. So we use SR560 to amplify it by a factor of 100 means 40dB. So the shot noise level should be -92dBV.
First, we measured the noise spectrum when one of the PD is blocked. The result is shown in the attached figure one. It is clear that not even part of noise spectrum is dominated by shot noise. Even at highest frequency, noise level is around -88dBV.
Then we decide to change pitch of steering mirror to change power on the second PD. Maybe this can bring beam away from optimal position. Or the change of beam direction reduce detection efficiency. In the end, we reduced common noise rejection from -30dBV to -74dBV, means we reduce the modulated signal by a factor of 100. The result is shown in the attached figure 2.
Finally, we turned off modulation. And then we look at the signal from homodyne on spectrum analyzer. The result is shown in the attached figure 3. Then we improve the locking condition of IRMC, make the locking point as close to peak as possible. Then the signal becomes figure 4. This is reasonable because around peak the amplitude change is less. Also seems the noise comes from amplitude noise.
To do list:
1.We can also change the pitch of BS to see if we can increase signal on the first PD.
2. Try to improve locking performance.
3. To see difference when noise eater is engaged and not.
4. Noise hunting

[Aritomi, Yuhang]
The procedure we did for aligning homodyne is:
1. Lock IMC and driving IMC end mirror with a ramp signal(1kHz 50mV). This will drive the output of IMC moving around the locking point. If we lock well on the peak value of IMC transmission. We will see the double frequency of the end mirror driving frequency. In this case we can use the modulated IMC transmission signal.
2. Connect homo-dyne and put it roughly in a good height and good horizental position. Here look at modulated IMC tra signal and make it maximum. Then fix homodyne.
3. Put lens in roughly the corresponding focal length of lens away from homodyne. Then roughly make the signal maximum again. Then fix lens.
4. Adjust lens so that we can further maximize.
After that we take the modulated signal both on oscilloscope and spetrum analyzer. Also take the signal when one of PD blocked or neither of them blocked.
The first three attached figure is for oscilloscope. The first one is when the second PD is blocked. The second figure is when the first PD is blocked. The third figure is when neither if them is blocked. (first PD is the nearest PD to IMC) (note here IMC is not locked on peak)
The next three figure is for block second PD, block first PD and no block. (here IMC is locked on peak)
Optimization is still needed.

[Aritomi, Yuhang]
Since we receive message from Henning, the voltage from pin 1 to 3 to 5 should be from -19 to 0 to 19V. So it means we should swap the connector. After do that we can see the signal on oscilloscope and its value is reasonable. Also the current coming out from power supply becomes not zero. As shown in the last attached figure. So we can say homodyne works well now.
After that, I measured again the dark noise of homo-dyne.
Firstly, I measured the electronic noise while SR560 is set gain as 1000(means 60dB). The result is shown in the first attached figure.
Then I measured the signal from homo-dyne whlie no light going inside homo-dyne. The result indicates the dark noise of homo-dyne is -93.2-60 = -153.2 dBV/rtHz. The result from Henning is -156.2dBVrms/rtHz(shown in attached figure 3). And since there is relationship Vrms = V/sqrt(2). This factor sqrt(2) is exactly 3dB difference. This means our measurement perfectly matches Henning's measurement.

The first two pictures shows the set-up of our power supply. As you can see from the picture, the voltage we are giving is +/- 19V.
The third picture shows the connection of the cable to the voltage supply. The connection follows the content of labels.
Finally, as pointed by Henning, I checked the voltage between pin 1, pin 5 and pin 3. These are shown in the attached picture 4 and 5. The voltage between pin 1 and 3 is 19V. While the voltage between pin 5 and 3 is -19V.
This result means the voltage going to homodyne should be fine.

[Aritomi, Eleonora, Matteo, Yuhang]
Investigations carried on yesterday and today seem to confirm that we might have a problem in the powering of the homodyne detector.
The main clue is that we cannot observe any change in its output signal when the power cable is connected and when it is not.
We do see a change in the spectrum of the signal when we shine a laser with an amplitude modulation (see Fig. 1) but the signal is the same even if we disconnect the power cable.
We checked the power cable connections and confirmed that the voltage arriving to the homodyne through it is the correct one.
Fig. 2, 3 show the homodyne detector and the setup.

All the electronics in the cleanbooth have been switched off, except for the LabView supervisor PC.
The oplev lasers, the coildrivers and the target labview PCs are also off, both in the central and in the end room.
All the used cleansuits, hoods and shoes have been sent to laundry.

Participaint: Aritomi, Eleonora, Matteo and Yuhang
First we checked the homodyne dark noise. The purpose is to compare this result with Henning's result. We are using network analyzer, but itself's noise level is not low enough to see the dark noise of homo-dyne. So we decided to use Stanford research SR560 to amplify the homodyne noise. We gave it a factor of 100(40dB). Then we got the result as shown in attached figure 1 and 2. The difference is they have a marker located at 1kHz and 10kHz, others are the same. We can see below 3kHz, the noise should be dominated by electronic noise. And above 3kHz, the noise level should be the noise level of homodyne. In this case, the real noise level of homodyne should be around -122-40 = -162dBV/sqrt(Hz). However, the result from Henning is
-156dBV/sqrt(Hz). This is not a neligible difference.
We are arriving to the point to operate homo-dyne. We power up homo-dyne with a DC voltage supply at +/- 19V. For checking both the alignment of the beam inside homodyne and also if the homodyne work well, we send an amplitude modulation to the main laser beam. At the beginning, we send a 100mV pk-pk and 1kHz signal into the current control of main laser. However, this modulation seems not stable. Then we decide to modulate the PZT of IRMC, in this case, we moved the locking point of IRMC around the TEM00 peak. Then the modulation became very clear. Here the clear or stable means if the signal we see on the oscilloscope shows clear line or can be triggered.
We also find out that the the even we send the voltage +/- 19V to homodyne. The current going inside seems like zero.
And when we see the 1kHz amplitude modulation signal on the monitor PD with amplitude of several volts. We can only see the same signal on homodyne with amplitude of several millivolts. Seems like we didn't switch on homodyne.
The last figure shows the homodyne on our bench.

Participaint: Aritomi, Eleonora and Yuhang
According to the optical layout, the lens should be 250mm before beam waist. And the beam waist is 390um. By putting a 30mm lens, we found the beam waist after lens will be 22um. The distance of the waist from the lens is 29mm. The corresponding rayleigh range is 1.4mm. The aperature of homodyne detector is 0.5mm, so we should make the going inside beam smaller than 100um. The range of beam smaller than 100um is 12mm.
The measurement result agrees with the simulation. The result is attached in the figure2.


The attached figure and PDF is the same. They show the space we may have the leg for breadboard. The space in this case is roughly 90cm*75cm. It covers all the space for PLL and two auxiliary laser heads.

1. The p-pol beam is very sensitive to the mirror mount. Even you touch the mirror mount, the alignmet condition will be changed.
2. The OPO's BAB transmission is also very sensitive to mirror mount.
3. The PLL locking is not robust enough. It can only last for several tens of minutes. And very diffcult to acquire, we can close and open software again to make this better.
4. The second Faraday's mount keeps moving. And its transmission has a lot of scattering.

Participaint: Eleonora and Yuhang
We convert BAB into p-pol and then measured the beat between it and the LO at the homodybne BS.
The result is shown in the attached picture. The visibility is (Vmax-Vmin)/(Vmax+Vmin) = (4.26-1.26)/(4.26+1.26) = 0.5435
The BAB power is 0.125mW and LO power is 1.24mW. The expected visibility is 2*sqrt(P1*P2)/(P1+P2) = 0.5768
So from these values, we should have 0.5435/0.5768 = 94.22% of matching of two beams. This is complaint with the matching we measured before.
There is one thing block us from measuring visibility, it is the use of lenses. Before we used lense, we can see the beam is flashing but cannot see this signal on oscilloscope.
By considering the level of visibility, I did the simulation of degradation. As shown in the attached figure two. Now because of only the reason of visibility, we degrade squeezing by 2.74dB.(I assume here the initial squeezing level is -10dB, the degradation is different for different initial squeezing level)

Participaint: Aritomi and Yuhang
We decide to change the position of OPO transmission PBS because of the space limitation. We also put the ''flip'' mirror on the translation stage. By moving this mirror, we will switch its direction to homodyne or to filter cavity. The scheme now is to use the two steering mirror before this ''filp'' mirror to do alignment for filter cavity. And adjust this ''flip'' mirror to align the beam into homodyne. Because this ''filp'' mirror has three adjustable knobs.
According to the design shown in the first attached figure, we aligned the BAB into AMC. However, the matching is not as good as LO.
From all the peak values, we can compute the matching is (0.51-0.0065)/((0.0115+0.0153+0.0173-0.0065*3)+0.51-0.0065) = 95.34%
We didn't see the visibility and found out the reason is we forgot to convert BAB into P-pol.

Participaint: Aritomi and Yuhang
Since we found PDA05CF2 (thorlabs InGAS PD) is more suitable for infrared signal. We decided to use it for OPO locking. Before we took the filter cavity locking PD for OPO locking because it has a DC channel. Since we have a better one, we decide to put this qubigPD (with DC) back for filter cavity locking.
Besides, according to simulation, the OPO reflection error signal is 5 times smaller than transmission. And we improve the error signal by a factor of 10. So in principle, now the error signal from OPO reflection should be also enough to lock OPO.

This PD has only one output channel. RF and DC components are separated by using a minicircuit Bias-Tee ZFBT-4R2G+ (datasheet uploaded on the wiki).
The RF part is sent to the mixer for PDH demodulation, the DC is used to monitor the lock status.

Actually we did this replacement long time ago. But I would like to put here some reference value of this PD. We should have error signal pk-pk value more than 120mV. And DC value more than 3.3V. Note here the impedence is .
Besides, we can also see that this PD has a better performace than the Qubig we had. There is a result of previous measurement of Qubig PD. Compared with that, we improve the error signal by a factor of 10.
Actually, there is only a factor of 3 or 4 difference between Qubig PD and thorlabs PD's resoponsivity.
This PD has only one output channel. RF and DC components are separated by using a minicircuit Bias-Tee ZFBT-4R2G+ (datasheet uploaded on the wiki).
The RF part is sent to the mixer for PDH demodulation, the DC is used to monitor the lock status.

Participaint: Aritomi and Yuhang
This beam splitter will be used for homodyne. And the balance of these two power value is very important. So we decide to measure this value.
We bought two BSW41-1064, according to specfication. The overall performance is T=50+/-5% while R=50+/-5%. So R/T should be between 0.905 and 1.105.
This power ratio depends also on polarization. According to thorlab website, p-pol should be closer to 50:50. We measured the power ratio relationship with polarization for one mirror. The result is shown in the attached figure. It seems there is relationship between polarization and power ratio. But it is small dependence. We can also see this mirror is still within the error range of power ratio provided by thorlab. But it is not as Matteo suggested.
Then we changed it to the second mirror. Then the power ratio becomes R:T=607.5/615.4=0.987. This is much better.

Matteo and Aritomi recovered the vacuum pump for filter cavity. The procedure is to first bring back rotative pump and after it reachs a certain level. Turn on the molecular pump. There are some valves to seperate pump. According to the sequence of the pump on, open them.

Manuel, Victor
1310nm probe
we aligned the pump beam to maximize the AC signal of the surface reference sample, and we made a calibration scan with pump power 35mW.
Then we increased the power at the maximum and measured the crystalline coating with three different probe power: DC=1.01, DC=1.62, DC=2.54.
Then we plot the three scans together and the result is that the AC/DC overlaps.
We also repeated a scan with higher resolution in z.

Participaint: Aritomi and Yuhang
We aligned the local oscillator into the alignent mode cleaner(AMC). The method is to remove input and output mirror and then make sure the beam going through MC through the center. And the beam should be flat horizentally and vertically. The reason why we aligned LO first is it is easier because it is much brighter than the OPO transmission. Then we can use this as a reference for the alignment of the second beam.
The resule of this match is shown in the first and second attached figure. The matching is (1.75-0.012-(0.02-0.012))/(1.75-0.012)=99.5%